# The Host Galaxies of Fast Radio Bursts Track a Combination of Stellar Mass and Star Formation, Similar to Type Ia Supernovae

### v1

@@ 4.2 Marginalized Distribution Functions
To further investigate the results discussed above, we plot in Figure 4 the univariate marginalized CDFs as a function of either stellar-mass (top panels) or SFR (bottom panels). To limit the effects of redshift evolution we separate these CDFs into three redshift bins selected as: $z<0.2$, $0.40.2<z<$ 0.4, and $z>0.4$. These particular redshift bins were chosen to align with those in K. Sharma et al. (2024) and N. Loudas et al. (2025) in order to better facilitate comparison with these works. In each panel we plot the CDFs of the mass-weighted model (yellow), the SFR-weighted model (bright red), the K. Sharma et al. (2024) SFR+Z model (dark red), and our preferred mixed model (green) with $A/B=10^{-10.5}\,{\rm yr}^{-1}$. We also plot the FRB host-galaxy data in blue along with the associated 95% confidence interval (CI) obtained by bootstrapping, using the following procedure. For the given dataset of size $N=51$ listing the FRB host-galaxy properties (either stellar-mass or SFR), we create a bootsrapped dataset by sampling the original dataset $N$ times, with repetitions allowed. We generate 1,000 bootstrapped samples in this way, and calculate the 95% CI of the resulting CDFs. Figure 4 also lists the p-values resulting from a KS test of each model against the FRB data, quoted in the top-left corner of each plot and following the color-coding convention for the models.
Figure 4: Marginalized stellar-mass (top panels) and SFR (bottom) distribution functions, split into three redshift bins. Blue curves show the CDFs of the FRB hosts within each redshift bin, and shaded grey areas show the 95% CI on these CDFs obtained by bootsrapping. The colored curves in each panel show different models discussed in the text: the mass-weighted model (yellow), the SFR-weighted model (bright red), the SFR+Z model proposed by K. Sharma et al. (2024) (dark red), and our preferred mixed model for which $A/B=10^{-10.5}\,{\rm yr}^{-1}$ (green). The p-values resulting from KS tests of each model against the FRB data are listed in the top-left corner of each plot. See §4.2 for further details. The model samples in each subplot are generated at a redshift corresponding to the median redshift of FRB host galaxies within that bin.
@@
Similarly, Figure 4 shows that the mass-weighted model systematically over-predicts the stellar-mass compared to the observed FRB host galaxies. This leads us to disfavor the mass-weighted model, even though it cannot be strictly ruled out by our primary analysis (it has a marginal p-value of $\approx 0.07$; see Table 1). As previously discussed, the methodology of our primary analysis is conservative in the sense that models that achieve a high p-value may not necessarily be representative of the underlying distribution of the data, however models that garner low p-values can robustly be ruled out. The fact that the mass-weighted model does not appear to fit the marginalized CDFs is therefore not in tension with its moderate p-value.
@@
We note that the low redshift distributions ($z\lesssim 0.2$) are subject to uncertainties associated with the choice of how to extend the J. Leja et al. (2022) distribution function (which is only defined at $30.2<z<$ 3). As discussed towards the end of §3.1, our currently-adopted approach may potentially overestimate the theoretical (model) SFRs at $z<0.2$ by as much as $\sim 0.3\,{\rm dex}$. If this is true, then the theoretical curves in the bottom-left panel of Figure 4 should shift leftwards. This would likely have the effect of improving agreement between our preferred mixed model and the FRB data. It may also similarly affect the SFR-weighted and SFR+Z models, but we leave this for future investigation.
@@ 5 Discussion and Summary
In this work we have presented a new framework for statistically testing models that describe the host-galaxy distributions of astrophysical sources. Our framework has the advantage of being applicable to sources that span a wide range in redshift, and accounts for redshift evolution in a way that is unbiased and model-agnostic, avoiding possible pitfalls to other approaches in the literature (such as binning or ‘transposition’; see §2). Equally-important, our methodology uses the multi-variate nature of the data in a self-consistent way—leveraging both host-galaxy stellar-mass and SFR measurements in a joint analysis, rather than examining each (or only one) property independently.
@@
  • Using our method we rule out the hypothesis that FRBs track SFR. We cannot strictly rule out the hypothesis that FRBs track stellar-mass, but we consider this scenario unlikely given the marginal p-values and the poor match between the resulting CDFs and the FRB data (Table 1; Figure 4).
@@
  • This preferred model is remarkably similar to inferences for Type Ia SNe, where an identical model with similar $A/B$ values works well (E. Scannapieco & L. Bildsten, 2005). This suggests that FRB progenitors and Type Ia SNe may form through similar channels, potentially implicating accretion-induced collapse as an FRB formation pathway (e.g., B. Margalit et al., 2019; K. Kremer et al., 2021; W. Lu et al., 2022; T. Eftekhari et al., 2025).
  • A prediction of this model is that the fraction of quiescent, transitioning, and star-forming host galaxies should be approximately 10%, 10%, and 80%, respectively (see §4.1). This is broadly consistent with the current sample of FRB hosts, although quiescent galaxies are underrepresented in the current sample. We therefore predict that more quiescent FRB hosts will be found. This is consistent with the recent discovery of FRB 20240209A (V. Shah et al. 2025; T. Eftekhari et al. 2025; which is not included in out data set). Another prediction is that the fraction of quiescent galaxies should decrease at higher redshifts.
@@
  • We also test the hypothesis that all FRBs are produced in GCs. Based on host-galaxy demographics alone, we rule out this hypothesis at the $p=0.02$ level. To the best of our knowledge, this is the first time that this scenario has been explicitly ruled out. Of course, this does not preclude the possibility that only a subset of FRBs originate from GCs, consistent with at least one source (F. Kirsten et al., 2022).
@@
Our finding that the host-galaxy distribution of FRBs is best modeled as a mixed distribution that follows both stellar mass and star formation can be interpreted in two ways: (i) that there are two distinct formation channels for FRB sources, one that tracks star formation and another that tracks stellar mass, or (ii) that FRB sources are born through a single channel that has an extended delay-time distribution. We suggest that the first interpretation can be tested by comparing the properties of FRBs that occur in galaxies whose ${\rm sSFR}>A/B\approx 0.03\,{\rm Gyr}^{-1}$ with FRBs in galaxies where ${\rm sSFR}<A/B$. This separates FRBs into two classes based on whether the source was more likely to have formed though the SFR- or stellar-mass-associated channels. If such sources are truly distinct, one might expect their radio properties to differ from one another (although this must not necessarily be the case). The second interpretation of our mixed model is consistent with the standard view of Type Ia SNe and short GRBs which are thought to form via channels that have extended time-delay distributions, such as compact object mergers. Indeed, our results are consistent with previously suggested similarities between the host-galaxies of FRBs, Type Ia SNe and short GRBs (e.g., B. Margalit et al., 2019; T. Eftekhari et al., 2025)
@@
The possible connection between FRBs and Type Ia SNe is also supported by direct comparison of FRB host-galaxy properties with Type Ia SNe hosts. A recent comparison of this type can be found in K. Sharma et al. (2024). Extended Data Figure 8 and Supplementary Table 5 in K. Sharma et al. (2024) show that the stellar-mass and SFR CDFs of Type Ia SNe hosts agree well with the current sample of FRB hosts. Similarly, Extended Data Figure 7 in that work shows that the offsets of Type Ia from their host galaxies agree well with FRB offsets. These facts further support our current findings, obtained using a completely different methodology and without directly relying on Type Ia SNe data.
@@
We encourage future work using our formalism to further investigate FRB progenitor models, especially as the sample of localized FRBs increases in size. For example, models that depend non-linearly on mass or SFR, or a broader class of metallicity-dependent models could be investigated. Furthermore, a more detailed treatment of the optical selection function, as recently proposed by N. Loudas et al. (2025), could be incorporated into our analysis. Finally, we note that our current framework requires extending the empirical galaxy distribution function at low redshifts $z\lesssim 0.2$. Such an extension was also necessary in the recent works by K. Sharma et al. (2024) and N. Loudas et al. (2025). In those works, the low-redshift extension extrapolates fitting formulae to the stellar-mass function and star-forming sequence obtained at higher redshifts. In our current work we instead assume that there is no redshift evolution below a cutoff redshift of $z=0.25$. These choices can potentially impact the results (see discussion towards end of sections 3.1 and 4.2), and we encourage future work to investigate these issues.
@@
The source code used in our study is publicly available at https://github.com/hoasaf3/host-galaxies-stats, and is designed to be applicable to broader contexts beyond the specific scope of FRB host-galaxies. We caution that it is important that application of this framework be limited to galaxies whose photometry and/or spectroscopy have been modeled using the non-parametric SED-modeling code Prospector (J. Leja et al., 2017; B. D. Johnson et al., 2021). This is important because our formalism uses a galaxy distribution that is trained on galaxies that have been modeled using this code, and recent work has shown that the choice of SED-modeling code can systematically affect the inferred galaxy properties (J. Leja et al., 2019, 2022).
We thank Alexa Gordon for helpful conversations and for commenting on an early version of this work, and Nick Loudas for insightful comments and for helping facilitate a comparison of our respective works. Author’s Note (AH): This paper has taken many turns since its inception, but its earliest draft bore the title “Can Babies Be Born in a Nursing Home?”—a nod to the question that drove the initial version of this work: how likely is it for an FRB, widely considered to originate from a young magnetar, to come from a GC—an environment dominated by old stars—as in the case of FRB 20200120E. The scope has since broadened, but I still wanted to give that title a small place in the final version. , This Appendix contains Tables listing several properties relevant to the FRB host-galaxy sample used in this work.

### v2

@@ 4.2 Marginalized Distribution Functions
To further investigate the results discussed above, we plot in Figure 4 the univariate marginalized CDFs as a function of either stellar-mass (top panels) or SFR (bottom panels). To limit the effects of redshift evolution we separate these CDFs into three redshift bins selected as: $z<0.2$, $0.2<z<0.4$, and $z>0.4$. These particular redshift bins were chosen to align with those in K. Sharma et al. (2024) and N. Loudas et al. (2025) in order to better facilitate comparison with these works. In each panel we plot the CDFs of the mass-weighted model (yellow), the SFR-weighted model (bright red), the K. Sharma et al. (2024) SFR+Z model (dark red), and our preferred mixed model (green) with $A/B=10^{-10.5}\,{\rm yr}^{-1}$. We also plot the FRB host-galaxy data in blue along with the associated 95% confidence interval (CI) obtained by bootstrapping, using the following procedure. For the given dataset of size $N=51$ listing the FRB host-galaxy properties (either stellar-mass or SFR), we create a bootsrapped dataset by sampling the original dataset $N$ times, with repetitions allowed. We generate 1,000 bootstrapped samples in this way, and calculate the 95% CI of the resulting CDFs. Figure 4 also lists the p-values resulting from a KS test of each model against the FRB data, quoted in the top-left corner of each plot and following the color-coding convention for the models.
Figure 4: Marginalized stellar-mass (top panels) and SFR (bottom) distribution functions, split into three redshift bins. Blue curves show the CDFs of the FRB hosts within each redshift bin, and shaded grey areas show the 95% CI on these CDFs obtained by bootsrapping. The colored curves in each panel show different models discussed in the text: the mass-weighted model (yellow), the SFR-weighted model (bright red), the SFR+Z model proposed by K. Sharma et al. (2024) (pink), and our preferred mixed model for which $A/B=10^{-10.5}\,{\rm yr}^{-1}$ (green). The p-values resulting from KS tests of each model against the FRB data are listed in the top-left corner of each plot. See §4.2 for further details. The model samples in each subplot are generated at a redshift corresponding to the median redshift of FRB host galaxies within that bin.
@@
Similarly, Figure 4 shows that the mass-weighted model systematically over-predicts the stellar-mass compared to the observed FRB host galaxies. This is not surprising, given that the mass-weighted model is marginally ruled out by our analysis (it has a p-value of $\approx 0.05$; see Table 1). As previously discussed, the methodology of our primary analysis is conservative in the sense that models that achieve a high p-value may not necessarily be representative of the underlying distribution of the data, however models that garner low p-values can robustly be ruled out. The fact that the mass-weighted model does not appear to fit the marginalized CDFs is therefore not in tension with its moderate p-value.
@@ 5 Discussion and Summary
In this work we have presented a new framework for statistically testing models that describe the host-galaxy distributions of astrophysical sources. Our framework has the advantage of being applicable to sources that span a wide range in redshift, and accounts for redshift evolution in a way that is unbiased and model-agnostic, avoiding possible pitfalls to other approaches in the literature (such as binning or ‘transposition’; see §2). In particular, our method avoids redshift-dependent selection effects by using the conditional probability $\rho\left(M,{\rm SFR}\,|z\right)$ and constructing redshift-matched mock catalogs (whereby each mock source is matched to the redshift of a real source in the astrophysical sample). Additionally, our methodology uses the multi-variate nature of the data in a self-consistent way—leveraging both host-galaxy stellar-mass and SFR measurements in a joint analysis, rather than examining each (or only one) property independently.
@@
  • Using our method we rule out the hypothesis that FRBs track SFR. We can only marginally rule out the hypothesis that FRBs track stellar-mass, but we consider this scenario unlikely given the marginal p-values and the poor match between the resulting CDFs and the FRB data (Table 1; Figure 4).
@@
  • This preferred model is similar to inferences for Type Ia SNe, where an identical model with similar $A/B$ values works well (E. Scannapieco & L. Bildsten, 2005). This suggests that FRB progenitors and Type Ia SNe may form through similar channels, potentially implicating accretion-induced collapse as an FRB formation pathway (e.g., B. Margalit et al., 2019; K. Kremer et al., 2021; W. Lu et al., 2022; T. Eftekhari et al., 2025).
  • A prediction of this model is that the fraction of quiescent, transitioning, and star-forming host galaxies should be approximately 10%, 10%, and 80%, respectively (see §4.1). This is broadly consistent with the current sample of FRB hosts, although quiescent galaxies are underrepresented in the current sample. We therefore predict that more quiescent FRB hosts will be found. This is consistent with the recent discovery of FRB 20240209A (V. Shah et al. 2025; T. Eftekhari et al. 2025; which is not included in our data set). Another prediction is that the fraction of quiescent galaxies should decrease at higher redshifts.
@@
  • We also test the hypothesis that all FRBs are produced in GCs. Based on host-galaxy demographics alone, we rule out this hypothesis at the $p=0.01$ level. Of course, this does not preclude the possibility that only a subset of FRBs originate from GCs, consistent with at least one source (F. Kirsten et al. 2022; see also A. Rao et al. 2025).
@@
Our finding that the host-galaxy distribution of FRBs is best modeled as a mixed distribution that follows both stellar mass and star formation can be interpreted in two ways: (i) that there are two distinct formation channels for FRB sources, one that tracks star formation and another that tracks stellar mass, or (ii) that FRB sources are born through a single channel that has an extended delay-time distribution. We suggest that the first interpretation can be tested by comparing the properties of FRBs that occur in galaxies whose ${\rm sSFR}>A/B\approx 0.03\,{\rm Gyr}^{-1}$ with FRBs in galaxies where ${\rm sSFR}<A/B$. This separates FRBs into two classes based on whether the source was more likely to have formed though the SFR- or stellar-mass-associated channels. If such sources are truly distinct, one might expect their radio properties to differ from one another (although this must not necessarily be the case). The second interpretation of our mixed model is consistent with the standard view of Type Ia SNe and short GRBs which are thought to form via channels that have extended time-delay distributions, such as compact object mergers. Indeed, our results are consistent with previously suggested similarities between the host-galaxies of FRBs, Type Ia SNe and short GRBs (e.g., B. Margalit et al., 2019; T. Eftekhari et al., 2025; J. H. Chen et al., 2024). Our conclusion—derived from FRB host galaxies—that FRB sources track a combination of both SFR and stellar-mass supports recent suggestions based on modeling the FRB redshift evolution and luminosity function (J.-f. Mo et al. 2025; O. Gupta et al. 2025; although see C. W. James et al. 2022; K. Shin et al. 2023 who, using similar methods, instead favor a scenario in which FRBs track SFR, or T. Hashimoto et al. 2022; J. H. Chen et al. 2024 who conversely find that FRBs more-closely track stellar-mass).
@@
It is important to note that there are various caveats to our analysis, and our findings should be interpreted with caution. In particular, there is a risk that the similarity in the inferred $A/B$ values for Type Ia SNe and FRBs is driven by a confounder, rather than reflecting an intrinsic similarity between the progenitors of these distinct transients. We note however that the possible connection between FRBs and Type Ia SNe is separately supported by direct comparison of FRB host-galaxy properties with Type Ia SNe hosts. A recent comparison of this type can be found in K. Sharma et al. (2024). Extended Data Figure 8 and Supplementary Table 5 in K. Sharma et al. (2024) show that the stellar-mass and SFR CDFs of Type Ia SNe hosts agree well with the current sample of FRB hosts. Similarly, Extended Data Figure 7 in that work shows that the offsets of Type Ia from their host galaxies agree well with FRB offsets. These facts further support our current findings, obtained using a completely different methodology and without directly relying on Type Ia SNe data.
@@
We encourage future work using our formalism to further investigate FRB progenitor models, especially as the sample of localized FRBs increases in size. For example, models that depend non-linearly on mass or SFR, or a broader class of metallicity-dependent models could be investigated. Furthermore, a more detailed treatment of the optical selection function, as recently proposed by N. Loudas et al. (2025), could be incorporated into our analysis. Finally, we note that our current framework requires extending the empirical galaxy distribution function at low redshifts $z\lesssim 0.2$. Such an extension was also necessary in the recent works by K. Sharma et al. (2024) and N. Loudas et al. (2025). Using this method, the low-redshift extension extrapolates fitting formulae to the stellar-mass function and star-forming sequence obtained at higher redshifts. A different approach, which we have tested as well, is to assume that the galaxy distribution evolves very little between $z=0.2$ and $z=0$, and therefore that the distribution function is constant at these redshifts. This approach yields very similar results, and does not alter our primary conclusions. Both choices have drawbacks, and the only self-consistent approach would be to train a model on field-galaxy data that goes down to $z=0$. In our current work we choose a redshift extension that is consistent with previous works (K. Sharma et al., 2024; N. Loudas et al., 2025). Although this method introduces some model dependence, it is more realistic than assuming a no redshift evolution (e.g., P. Madau & M. Dickinson, 2014).
@@
The source code used in our study is publicly available [^5] https://github.com/hoasaf3/host-galaxies-stats (A. Horowicz, 2025), and is designed to be applicable to broader contexts beyond the specific scope of FRB host-galaxies. We caution that it is important that application of this framework be limited to galaxies whose photometry and/or spectroscopy have been modeled using the non-parametric SED-modeling code Prospector (J. Leja et al., 2017; B. D. Johnson et al., 2021). This is important because our formalism uses a galaxy distribution that is trained on galaxies that have been modeled using this code, and recent work has shown that the choice of SED-modeling code can systematically affect the inferred galaxy properties (J. Leja et al., 2019, 2022).
We thank Alexa Gordon for helpful conversations and for commenting on an early version of this work, and Nick Loudas for insightful comments and for helping facilitate a comparison of our respective works. We also thank the anonymous referee for comments and suggestions that helped improve this manuscript. Author’s Note (AH): This paper has taken many turns since its inception, but its earliest draft bore the title “Can Babies Be Born in a Nursing Home?”—a nod to the question that drove the initial version of this work: how likely is it for an FRB, widely considered to originate from a young magnetar, to come from a GC—an environment dominated by old stars—as in the case of FRB 20200120E. The scope has since broadened, but I still wanted to give that title a small place in the final version. This Appendix contains Tables listing several properties relevant to the FRB host-galaxy sample used in this work.